Consider a mass spring suspended from a rigid support. Let v be the speed imparted to the mass , R be the length  of the string . If v = (4gR)½ , at what angle with the horizontal will the string be slack ?

Grade:11

1 Answers

Bhavasagar
6 years ago
Given,v=√4gR. If ∆ is the angle made by string with horizontal when it is slack. From FBD of mass. tan∆=weight/Centripetal force. tan∆=mg/mv^2/R. tan∆=mgR/m4gR. tan∆=1/4. ∆=tan^-1(1/4)